OpenStudy (anonymous):

Polar coordinates of a point are given. Find the rectangular coordinates of the point. (-5, -180°) A. (-5, 0) B. (0, -5) C. (0, 5) D. (5, 0) Reset Selection

OpenStudy (jdoe0001):

\(\bf (-5,-180^o)\qquad r=-5\qquad \theta=-180^o\\ \quad \\ x=rcos(\theta)\implies x=-5\cdot cos(-180^o)\\ \quad \\y=rsin(\theta)\implies y=-5\cdot sin(-180^o)\)

OpenStudy (jdoe0001):

get the values, plug them in at (x, y) for the rectangular form